OpenStudy (calculusxy):

Medal will be given! During their first home game, the Eagles soccer team drew a total of 132 fans. By the second game, this number grew to 162 fans. What was the percent increase in attendance between the first two home games of the season? Round your answer to the nearest tenth, if necessary. A. 30.3% B. 30% C. 22.7% D. 18.5%

OpenStudy (anonymous):

increase in the number of fans = 162 -132 = 30 Percentage increase is calculated with respect to the original quantity. \[% increase = \frac{ 30 }{ 132 } \times 100\]

OpenStudy (anonymous):

\[gain = \frac{ 30 }{ 132 } \times 100\]